OpenStudy (anonymous):

for what value of n does 10! divided by (7! times 3!) =n ?

hartnn (hartnn):

use the property that \(n! = n(n-1)!\) \((n-1)!=(n-1)(n-2)!\) and so on, like 4! = 4*3! = 4*3*2! =....

hartnn (hartnn):

so, what would you express 10! as ?

OpenStudy (anonymous):

10! =10 times 9 ???

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

i think that ! means to multiply down like 10x9x8x7x6x5x4x3x2x1

hartnn (hartnn):

10! = 10*9! but thats no use to us, we have 7! in the denominator, so, how about this, 10! = 10*9! = 10*9*8! = 10*9*8* 7! got this ?

OpenStudy (anonymous):

Thanks:)

hartnn (hartnn):

welcome :) and could you find n now ?